OpenStudy (anonymous):

Find the sum of a 10-term geometric sequence when the first term is 3 and the last term is 59,049 and select the correct answer below. pleeaaassseeee someone help

OpenStudy (michele_laino):

we have to apply the subsequent formula: \[\Large {a_{10}} = {a_1}{q^9}\] from which we have: \[\Large q = \sqrt[9]{{\frac{{{a_{10}}}}{{{a_1}}}}}\] where q is the constant of the geometric sequence. What is the value of q?

OpenStudy (michele_laino):

hint: \[\Large q = \sqrt[9]{{\frac{{{a_{10}}}}{{{a_1}}}}} = \sqrt[9]{{\frac{{59049}}{3}}} = ...?\]

OpenStudy (michele_laino):

the requested sum S, is given by the subsequent formula: \[\Large S = {a_1}\frac{{1 - {q^9}}}{{1 - q}} = ...?\] where a_1=3, namely is the first term of your geometric sequence

OpenStudy (anonymous):

note that once you find \(q\) you don't have to compute \(q^n\) again and you can just determine the sum using $$S=\frac{a_{10}-a_1}{q-1}$$

OpenStudy (condor34):

so whats the answer?